AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 292     Entry time: Sun Feb 27 01:46:56 2022
Author: David, Pierre-Michell 
Subject: Current status 

So far looks like with a temperature of 95K the target is maintained with a reasonable density throughout long periods of time.

Ring current 6.41e7 particles/s

Merger, Tape server (WS3): 1472 kB/s writing fine

Leakage current (WS4): 3.26 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): everything fine

Attachment 1: 2022-02-27_02-50-15exrate.png  154 kB  | Hide | Hide all
2022-02-27_02-50-15exrate.png
Attachment 2: 2022-02-27_02-49-47extot.png  164 kB  | Hide | Hide all
2022-02-27_02-49-47extot.png
Attachment 3: 2022-02-27_02-49-28xytot.png  180 kB  | Hide | Hide all
2022-02-27_02-49-28xytot.png
ELOG V3.1.4-unknown